My ages are: 10,50,26. How do I get those to add up to 150? What's the equation if it is even possible?
Found 2 solutions by jim_thompson5910, Alan3354:
Answer by jim_thompson5910(35256)

Currently they add up to 10+50+26 = 86

If you add x years to each age, then you get 10+x, 50+x, and 26+x

Add up the ages and force that result to be 150. Then solve for x.

(10+x)+(50+x)+(26+x) = 150

10+x+50+x+26+x = 150

3x+86 = 150

3x = 150-86

3x = 64

x = 64/3

x = 21.333

So there isn't a whole number of years that can pass by where your three ages add to 150.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by Alan3354(69443)

My ages are: 10,50,26. How do I get those to add up to 150? What's the equation if it is even possible?
----------------
How can you be 3 ages?
They add to 86, not 150.